OpenStudy (anonymous):

Find the following. assume that x can represent a real number. use absolute value notation when necessary sqrt16x^2 =

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{16x^2}\] \[\large \sqrt{(4x)^2}\] \[\large |4x|\] \[\large |4||x|\] \[\large 4|x|\] So \[\large \sqrt{16x^2}=4|x|\]

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

simplify by factoring 3sqrt 243^8 =

jimthompson5910 (jim_thompson5910):

Just remember that \[\large \sqrt{x^2}=|x|\] where 'x' is any real number

OpenStudy (anonymous):

thanks I got it

jimthompson5910 (jim_thompson5910):

Is that a cube root?